A new born baby after birth is thrown upward by her mother in the usual way (after birth child exercise) with an initial velocity of 10m/s from a height of 10 metres above the ground. Taken that the mother's hands are just at the 10metres height of throw, determine the total distance the child covered. ( Take g =10m/s ^2)
1
Expert's answer
2019-10-17T11:12:28-0400

A baby thrown upward with initial velocity 10 m/s will reach height


"h=\frac{v^2}{2g}=\frac{10^2}{2\cdot10}=5\text{ m}."


Then the baby will fall down to mother's hands the same 5 meters. The total distance the child will cover is 10 meters. It does not depend on the height of the mother above the ground.